OpenStudy (studybuddy22):

Part 1 [2 points] Is the following equation true? Part 2 [6 points] Use complete sentences to explain the properties used in making your decision. 3 log_2 x + 1/2 log_2 y - 3 log_2 z = log_2 x^3 sqrt(y)/ z^3

OpenStudy (mertsj):

\[3\log_{2}x+\frac{1}{2}\log_{2}y-3\log_{2}z=\log_{2}x ^{3}\frac{\sqrt{y}}{z ^{3}} \]